# Open dynamical systems with a moving hole

### v1

@@ 4. Intermediate value property for dimensions of $K^{\omega}$
$x^{m}-bx^{m-1}+1=0\quad\text{and}\quad x^{m}-(b-1)(x^{m-1}+x^{m-2}+\ldots+x+1)=0,$
@@
Let $\xi$ and $\zeta$ be totally distinct and progressively overlapping sequences in $(D_{b}^{m})^{\mathbb{N}}$, respectively. By Proposition 3.5, it suffices to show that
@@
Note that $\eta$ satisfies the equation $x^{m}-bx^{m-1}+1=0$, i.e., $b-\eta^{-(m-1)}=\eta$. Assume that for some $k\in\mathbb{N}$ we have $|\Sigma_{j+1}^{\omega}|-|\Sigma_{j+1}^{\xi}|\geq\eta(|\Sigma_{j}^{\omega}|-|\Sigma_{j}^{\xi}|)$ for all $11\leq j<k+m-$ 1. Then by Lemma 3.2 we have
@@
Take $\omega\in(D_{b}^{m})^{\mathbb{N}}$. Recall from (2.2) that for $\mathbf{c}=c_{1}c_{2}\ldots c_{k}\in\Sigma_{k}^{\omega}$ and $n\in\mathbb{N}$ we define
@@
For $k\geq m$ and $n\geq m$, by (2.6) and (2.7) in the proof of Lemma 2.2 we have showed that $|\Sigma^{\tau}_{n-m+1}|\leq|\Sigma_{k+n}^{\omega}(\mathbf{c})|\leq b^{m-1}|\Sigma^{\tau}_{n-m+1}|$ for some $\tau\in(D_{b}^{m})^{\mathbb{N}}$. By Lemma 4.1, for $k\geq m$ and $n\geq m$ we have
@@
where $N_{r}(F)$ is the smallest number of balls of diameter $r$ required to cover a set $F$.
@@
For any $\omega\in(D_{b}^{m})^{\mathbb{N}}$ we have
@@
$\ell_{n}:=\sum_{i=0}^{n}(p_{i}+q_{i})+mn\quad\text{for}\leavevmode\nobreak\ n\in\mathbb{N}_{0}.$
Let $L\big{(}\{p_{n}\},\{q_{n}\}\big{)}$ be the set of sequences $\omega\in(D_{b}^{m})^{\mathbb{N}}$ that is progressive overlapping at position-$k$ for all $\ell_{n}<k\leq\ell_{n}+p_{n+1}$ and is totally distinct at position-$k$ for all $\ell_{n}+p_{n+1}<k\leq\ell_{n}+p_{n+1}+q_{n+1}$. For $\omega\in(D_{b}^{m})^{\mathbb{N}}$ and $k\in\mathbb{N}_{0}$ write
@@
$\mathbf{N}^{\omega}_{k}$ $=\mathbf{N}^{\omega}_{k-1}A\quad\text{if $\omega$ is progressive overlapping at position-$k$},$
@@
$\mathbf{N}^{\omega}_{k}$ $=\mathbf{N}^{\omega}_{\ell_{n}}A^{k-\ell_{n}}\quad\text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ \ell_{n}<k\leq\ell_{n}+p_{n+1},$
$\mathbf{N}^{\omega}_{k}$ $=\mathbf{N}^{\omega}_{\ell_{n}+p_{n+1}}B^{k-\ell_{n}-p_{n+1}}\quad\text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ \ell_{n}+p_{n+1}<k\leq\ell_{n}+p_{n+1}+q_{n+1}.$
@@
(4.5) $C_{1}^{-1}\lambda^{k-\ell_{n}}\|\mathbf{N}^{\omega}_{\ell_{n}}\|\leq\|\mathbf{N}^{\omega}_{k}\|\leq C_{1}\lambda^{k-\ell_{n}}\|\mathbf{N}^{\omega}_{\ell_{n}}\|\quad\text{for}\leavevmode\nobreak\ \leavevmode\nobreak\ \ell_{n}<k\leq\ell_{n}+p_{n+1};$
@@
Write ~$\widetilde{\ell}_{n}:=\sum_{i=1}^{n}(p_{i}+q_{i})$. Note that $\ell_{n}=\sum_{i=1}^{n}(p_{i}+q_{i})+mn.$ Then (4.9) implies that ~$\widetilde{\ell}_{n}/\ell_{n}\to 1$ as $n\to\infty$.
@@
$\geq\frac{(\sum_{i=0}^{n}q_{i})\log\eta+(\sum_{i=0}^{n}p_{i})\log\lambda}{\widetilde{\ell}_{n}\log b}-\frac{n\log C+mn\log\lambda}{\widetilde{\ell}_{n}\log b}\rightarrow\alpha\quad\textrm{as}\leavevmode\nobreak\ n\to\infty;$
@@
$\leq\frac{(\sum_{i=0}^{n}q_{i})\log\eta+(\sum_{i=0}^{n+1}p_{i})\log\lambda}{(\widetilde{\ell}_{n}+p_{n+1})\log b}+\frac{n\log C}{\widetilde{\ell}_{n}\log b}\rightarrow\beta\quad\textrm{as}\leavevmode\nobreak\ n\to\infty.$
@@
$\geq\frac{(\sum_{i=0}^{n+1}q_{i})\log\eta+(\sum_{i=0}^{n+1}p_{i})\log\lambda}{\widetilde{\ell}_{n+1}\log b}-\frac{n\log C+mn\log\lambda}{\widetilde{\ell}_{n}\log b}\;\rightarrow\alpha\quad\textrm{as}\leavevmode\nobreak\ n\to\infty;$
@@
$\leq\frac{(\sum_{i=0}^{n}q_{i})\log\eta+(\sum_{i=0}^{n+1}p_{i})\log\lambda}{(\widetilde{\ell}_{n}+p_{n+1})\log b}+\frac{n\log C}{\widetilde{\ell}_{n}\log b}\rightarrow\beta\quad\textrm{as}\leavevmode\nobreak\ n\to\infty.$
@@
Let $\tau=\tau^{0}\tau^{1}\ldots=\omega^{\ell_{n}+m-1}\omega^{\ell_{n}+m}\ldots\in(D_{b}^{m})^{\mathbb{N}}$. Since $\omega\in L\big{(}\{p_{n}\},\{q_{n}\}\big{)}$, the sequence $\tau$ is progressively overlapping at position-$k$ for all $11\leq k\leq p_{n+1}-m+$ 1. By (2.6) in the proof of Lemma 2.2 and Remark 3.6, we have
@@
Let $\tau=\tau^{0}\tau^{1}\ldots=\omega^{k_{1}+m-1}\omega^{k_{1}+m}\ldots\in(D_{b}^{m})^{\mathbb{N}}$. Since $\omega\in L\big{(}\{p_{n}\},\{q_{n}\}\big{)}$, the sequence $\tau$ is totally distinct at position-$k$ for all $11\leq k\leq q_{n+1}-m+$ 1. By (2.7) in the proof of Lemma 2.2 and Remark 3.6, we have
@@
Fix any sequence $d_{0}d_{1}d_{2}\ldots\in D_{b}^{\mathbb{N}}$, and we construct a sequence $\omega=\omega^{0}\omega^{1}\ldots\in L\big{(}\{p_{n}\},\{q_{n}\}\big{)}$ as follows. We first set $d_{-(m-1)}=\ldots=d_{-2}=d_{-1}=0$. For $n\in\mathbb{N}_{0}$ and $0\leq k\leq p_{1}$ or $\ell_{n}+p_{n+1}+q_{n+1}<k\leq\ell_{n+1}+p_{n+2}$, let $\omega^{k}=d_{k-(m-1)}\ldots d_{k-1}d_{k}$. Then the sequence $\omega$ is progressively overlapping at position-$k$ for all $\ell_{n}<k\leq\ell_{n}+p_{n+1}$. For $\ell_{n}+p_{n+1}<k\leq\ell_{n}+p_{n+1}+q_{n+1}$ we define $\omega^{k}=\omega_{1}^{k}\omega_{2}^{k}\ldots\omega_{m}^{k}$ recursively. Assume that $\omega^{i}$ have been defined for all $0\leq i<k$. Then we define $\omega^{k}$ by letting $\omega^{k}_{m}=d_{k}$ and choosing $\omega^{k}_{j}\in D_{b}\setminus\{\omega^{k-m+j}_{m}\}$ for all $11\leq j\leq m-$ 1. So the sequence $\omega$ is totally distinct at position-$k$ for all $\ell_{n}+p_{n+1}<k\leq\ell_{n}+p_{n+1}+q_{n+1}$.
@@
$\widetilde{\varrho}(\mathbf{d},\mathbf{d}^{\prime})=b^{-\inf\{k\in\mathbb{N}_{0}:d_{k}\neq d_{k}^{\prime}\}}\quad\text{for }\leavevmode\nobreak\ \mathbf{d}=d_{0}d_{1}d_{2}\ldots,\leavevmode\nobreak\ \mathbf{d}^{\prime}=d_{0}^{\prime}d_{1}^{\prime}d_{2}^{\prime}\ldots\in D_{b}^{\mathbb{N}}.$
Then one can check that $\dim_{H}D_{b}^{\mathbb{N}}=1$ under the metric $\widetilde{\varrho}$. Recall the metric $\varrho$ on $(D_{b}^{m})^{\mathbb{N}}$ defined in (1.5). We have
@@
Finally we prove (iii). In the case that $10<s<t<$ 1, the sequences $\{p_{n}\}_{n=1}^{\infty}$ and $\{q_{n}\}_{n=1}^{\infty}$ constructed above satisfy $p_{n}\to\infty$ and $q_{n}\to\infty$. By Proposition 4.6, we have for any $\omega\in L\big{(}\{p_{n}\},\{q_{n}\}\big{)}$,
@@
5. Applications in badly approximable numbers, non-recurrence points and finiteness property
In this section we will prove Theorems 1.8 and 1.10 for the set of badly approximable numbers and non-recurrence points, respectively, and prove Theorem 1.12 that the finiteness property holds in $\left\{A_{\mathbf{d}}:\mathbf{d}\in D_{b}^{m}\right\}$.

### v2

@@ 4. Intermediate value property for dimensions of $K^{\omega}$
$x^{m}-bx^{m-1}+1=0\quad\text{and}\quad x^{m}-(b-1)(x^{m-1}+x^{m-2}+\cdots+x+1)=0,$
@@
Let $\xi$ and $\zeta$ be totally distinct and progressively overlapping sequences in $(D_{b}^{m})^{\mathbb{N}_{0}}$, respectively. By Proposition 3.5, it suffices to show that
@@
Note that $\eta$ satisfies the equation $x^{m}-bx^{m-1}+1=0$, i.e., $b-\eta^{-(m-1)}=\eta$. Assume that for some $k\in\mathbb{N}$ we have $|\Sigma_{j+1}^{\omega}|-|\Sigma_{j+1}^{\xi}|\geq\eta(|\Sigma_{j}^{\omega}|-|\Sigma_{j}^{\xi}|)$ for all $1\leq j<k+m-1$. Then by Lemma 3.2 we have
@@
Take $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$. Recall from (2.2) that for $\mathbf{c}=c_{1}c_{2}\ldots c_{k}\in\Sigma_{k}^{\omega}$ and $n\in\mathbb{N}$ we define
@@
For $k\geq m$ and $n\geq m$, by (2.6) and (2.7) in the proof of Lemma 2.2 we have showed that $|\Sigma^{\tau}_{n-m+1}|\leq|\Sigma_{k+n}^{\omega}(\mathbf{c})|\leq b^{m-1}|\Sigma^{\tau}_{n-m+1}|$ for some $\tau\in(D_{b}^{m})^{\mathbb{N}_{0}}$. By Lemma 4.1, for $k\geq m$ and $n\geq m$ we have
@@
where $N_{r}(F)$ is the smallest number of balls of diameter $r$ required to cover a bounded set $F$.
@@
For any $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$ we have
@@
$\ell_{n}:=\sum_{i=0}^{n}(p_{i}+q_{i})+mn\quad\text{for}\penalty 10000\ n\in\mathbb{N}_{0}.$
Let $L\big(\{p_{n}\},\{q_{n}\}\big)$ be the set of sequences $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$ that is progressively overlapping at position-$k$ for all $\ell_{n}<k\leq\ell_{n}+p_{n+1}$ and is totally distinct at position-$k$ for all $\ell_{n}+p_{n+1}<k\leq\ell_{n}+p_{n+1}+q_{n+1}$. For $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$ and $k\in\mathbb{N}_{0}$ write
@@
$\mathbf{N}^{\omega}_{k}$ $=\mathbf{N}^{\omega}_{k-1}A\quad\text{if $\omega$ is progressively overlapping at position-$k$},$
@@
$\mathbf{N}^{\omega}_{k}$ $=\mathbf{N}^{\omega}_{\ell_{n}}A^{k-\ell_{n}}\quad\text{for}\penalty 10000\ \penalty 10000\ \ell_{n}<k\leq\ell_{n}+p_{n+1},$
$\mathbf{N}^{\omega}_{k}$ $=\mathbf{N}^{\omega}_{\ell_{n}+p_{n+1}}B^{k-\ell_{n}-p_{n+1}}\quad\text{for}\penalty 10000\ \penalty 10000\ \ell_{n}+p_{n+1}<k\leq\ell_{n}+p_{n+1}+q_{n+1}.$
@@
(4.5) $C_{1}^{-1}\lambda^{k-\ell_{n}}\|\mathbf{N}^{\omega}_{\ell_{n}}\|\leq\|\mathbf{N}^{\omega}_{k}\|\leq C_{1}\lambda^{k-\ell_{n}}\|\mathbf{N}^{\omega}_{\ell_{n}}\|\quad\text{for}\penalty 10000\ \penalty 10000\ \ell_{n}<k\leq\ell_{n}+p_{n+1};$
@@
Write $\widetilde{\ell}_{n}:=\sum_{i=1}^{n}(p_{i}+q_{i})$. Note that $\ell_{n}=\sum_{i=1}^{n}(p_{i}+q_{i})+mn.$ Then (4.9) implies that $\widetilde{\ell}_{n}/\ell_{n}\to 1$ as $n\to\infty$.
@@
$\geq\frac{(\sum_{i=0}^{n}q_{i})\log\eta+(\sum_{i=0}^{n}p_{i})\log\lambda}{\widetilde{\ell}_{n}\log b}-\frac{n\log C+mn\log\lambda}{\widetilde{\ell}_{n}\log b}\rightarrow\alpha\quad\textrm{as}\penalty 10000\ n\to\infty;$
@@
$\leq\frac{(\sum_{i=0}^{n}q_{i})\log\eta+(\sum_{i=0}^{n+1}p_{i})\log\lambda}{(\widetilde{\ell}_{n}+p_{n+1})\log b}+\frac{n\log C}{\widetilde{\ell}_{n}\log b}\rightarrow\beta\quad\textrm{as}\penalty 10000\ n\to\infty.$
@@
$\geq\frac{(\sum_{i=0}^{n+1}q_{i})\log\eta+(\sum_{i=0}^{n+1}p_{i})\log\lambda}{\widetilde{\ell}_{n+1}\log b}-\frac{n\log C+mn\log\lambda}{\widetilde{\ell}_{n}\log b}\;\rightarrow\alpha\quad\textrm{as}\penalty 10000\ n\to\infty;$
@@
$\leq\frac{(\sum_{i=0}^{n}q_{i})\log\eta+(\sum_{i=0}^{n+1}p_{i})\log\lambda}{(\widetilde{\ell}_{n}+p_{n+1})\log b}+\frac{n\log C}{\widetilde{\ell}_{n}\log b}\rightarrow\beta\quad\textrm{as}\penalty 10000\ n\to\infty.$
@@
Let $\tau=\tau^{0}\tau^{1}\ldots=\omega^{\ell_{n}+m-1}\omega^{\ell_{n}+m}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$. Since $\omega\in L\big(\{p_{n}\},\{q_{n}\}\big)$, the sequence $\tau$ is progressively overlapping at position-$k$ for all $1\leq k\leq p_{n+1}-m+1$. By (2.6) in the proof of Lemma 2.2 and Remark 3.6, we have
@@
Let $\tau=\tau^{0}\tau^{1}\ldots=\omega^{k_{1}+m-1}\omega^{k_{1}+m}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$. Since $\omega\in L\big(\{p_{n}\},\{q_{n}\}\big)$, the sequence $\tau$ is totally distinct at position-$k$ for all $1\leq k\leq q_{n+1}-m+1$. By (2.7) in the proof of Lemma 2.2 and Remark 3.6, we have
@@
Fix any sequence $d_{0}d_{1}d_{2}\ldots\in D_{b}^{\mathbb{N}}$, and we construct a sequence $\omega=\omega^{0}\omega^{1}\ldots\in L\big(\{p_{n}\},\{q_{n}\}\big)$ as follows. We first set $d_{-(m-1)}=\cdots=d_{-2}=d_{-1}=0$. For $n\in\mathbb{N}_{0}$ and $0\leq k\leq p_{1}$ or $\ell_{n}+p_{n+1}+q_{n+1}<k\leq\ell_{n+1}+p_{n+2}$, let $\omega^{k}=d_{k-(m-1)}\ldots d_{k-1}d_{k}$. Then the sequence $\omega$ is progressively overlapping at position-$k$ for all $\ell_{n}<k\leq\ell_{n}+p_{n+1}$. For $\ell_{n}+p_{n+1}<k\leq\ell_{n}+p_{n+1}+q_{n+1}$ we define $\omega^{k}=\omega_{1}^{k}\omega_{2}^{k}\ldots\omega_{m}^{k}$ recursively. Assume that $\omega^{i}$ have been defined for all $0\leq i<k$. Then we define $\omega^{k}$ by letting $\omega^{k}_{m}=d_{k}$ and choosing $\omega^{k}_{j}\in D_{b}\setminus\{\omega^{k-m+j}_{m}\}$ for all $1\leq j\leq m-1$. So the sequence $\omega$ is totally distinct at position-$k$ for all $\ell_{n}+p_{n+1}<k\leq\ell_{n}+p_{n+1}+q_{n+1}$.
@@
$\widetilde{\varrho}(\mathbf{d},\mathbf{d}^{\prime})=b^{-\inf\{k\in\mathbb{N}_{0}:d_{k}\neq d_{k}^{\prime}\}}\quad\text{for }\penalty 10000\ \mathbf{d}=d_{0}d_{1}d_{2}\ldots,\penalty 10000\ \mathbf{d}^{\prime}=d_{0}^{\prime}d_{1}^{\prime}d_{2}^{\prime}\ldots\in D_{b}^{\mathbb{N}}.$
Then one can check that $\dim_{H}D_{b}^{\mathbb{N}}=1$ under the metric $\widetilde{\varrho}$. Recall the metric $\varrho$ on $(D_{b}^{m})^{\mathbb{N}_{0}}$ defined in (1.5). We have
@@
Finally we prove (iii). In the case that $0<s<t<1$, the sequences $\{p_{n}\}_{n=1}^{\infty}$ and $\{q_{n}\}_{n=1}^{\infty}$ constructed above satisfy $p_{n}\to\infty$ and $q_{n}\to\infty$. By Proposition 4.6, we have for any $\omega\in L\big(\{p_{n}\},\{q_{n}\}\big)$,
@@
5. Regularity of $K^{\omega}$
In this section, we investigate some special sequences $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$ for which $\dim_{H}K^{\omega}=\dim_{P}K^{\omega}$. First, we generalize progressively overlapping and totally distinct sequences. For $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$, define
[…]
Recall the matrices $A$ and $B$ in (3.10), and the numbers $\lambda$ and $\eta$ defined in Lemma 3.3. We also use the same notation $\mathbf{N}^{\omega}_{k}=(|\Sigma^{\omega}_{k+m}|,|\Sigma^{\omega}_{k+m-1}|,\ldots,|\Sigma^{\omega}_{k+1}|)$ for $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$ and $k\in\mathbb{N}_{0}$.
Proposition 5.1.
Let $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$.
(i) If
(5.1) $\lim_{n\to\infty}\frac{\#\big(\Lambda_{\mathrm{PO}}^{\omega}\cap(0,n]\big)}{n}=1,$
then we have
$\dim_{H}K^{\omega}=\dim_{P}K^{\omega}=\dim_{A}K^{\omega}=\frac{\log\lambda}{\log b}.$
(ii) If
$\lim_{n\to\infty}\frac{\#\big(\Lambda_{\mathrm{TD}}^{\omega}\cap(0,n]\big)}{n}=1,$
then we have
$\dim_{L}K^{\omega}=\dim_{H}K^{\omega}=\dim_{P}K^{\omega}=\frac{\log\eta}{\log b}.$
Before proving the proposition we remark that in (i) we may not have $\dim_{L}K^{\omega}=\frac{\log\lambda}{\log b}$. Note that one can easily construct a sequence $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$ such that (5.1) holds and the length of consecutive integers in $\Lambda^{\omega}_{TD}$ is arbitrarily long. Then by the same argument as in the proof of Proposition 4.6 (ii) we have $\dim_{L}K^{\omega}=\frac{\log\eta}{\log b}<\frac{\log\lambda}{\log b}$. Similarly, in (ii) the equality $\dim_{A}K^{\omega}=\frac{\log\eta}{\log b}$ may not hold.
Proof.
(i) If $\mathbb{N}\setminus\Lambda_{\mathrm{PO}}^{\omega}$ is a finite set, then the sequence $\omega$ is eventually progressively overlapping. By Remark 1.5, the set $K^{\omega}$ is $\frac{\log\lambda}{\log b}$-Ahlfors regular. In the following, we assume that the set $\mathbb{N}\setminus\Lambda_{\mathrm{PO}}^{\omega}$ is infinite. Write $\mathbb{N}\setminus\Lambda_{\mathrm{PO}}^{\omega}=\big\{n_{1},n_{2},n_{3},\ldots\big\}$ where $n_{1}<n_{2}<n_{3}<\cdots$. Let $n_{0}=0$. Note that $\omega$ is progressively overlapping at position-$k$ for any $n_{j}<k<n_{j+1}$. It follows from Lemma 3.2 that
$\mathbf{N}^{\omega}_{k}=\mathbf{N}^{\omega}_{n_{j}}A^{k-n_{j}}\quad\forall n_{j}<k<n_{j+1}.$
For $n_{j}\leq k<n_{j+1}$, by Lemma 3.4 we have
$\|\mathbf{N}^{\omega}_{k}\|$ $\geq C_{1}^{-1}\lambda^{k-n_{j}}\|\mathbf{N}^{\omega}_{n_{j}}\|\geq C_{1}^{-1}\lambda^{k-n_{j}}\|\mathbf{N}^{\omega}_{n_{j}-1}\|$
$\geq C_{1}^{-2}\lambda^{k-n_{j-1}-1}\|\mathbf{N}^{\omega}_{n_{j-1}}\|\geq C_{1}^{-2}\lambda^{k-n_{j-1}-1}\|\mathbf{N}^{\omega}_{n_{j-1}-1}\|$
$\geq\cdots\cdots$
$\geq C_{1}^{-j-1}\lambda^{k-j}\|\mathbf{N}^{\omega}_{0}\|.$
That is,
$\frac{\|\mathbf{N}^{\omega}_{k}\|}{\|\mathbf{N}^{\omega}_{0}\|}\geq\frac{\lambda^{\#(\Lambda_{\mathrm{PO}}^{\omega}\cap(0,k])}}{C_{1}^{1+\#((0,k]\setminus\Lambda_{\mathrm{PO}}^{\omega})}}\quad\forall k\in\mathbb{N}.$
By Theorem 1.1 and (5.1), it follows that
$\dim_{H}K^{\omega}\geq\frac{\log\lambda}{\log b}.$
By Proposition 4.2, we conclude that
$\dim_{H}K^{\omega}=\dim_{P}K^{\omega}=\dim_{A}K^{\omega}=\frac{\log\lambda}{\log b}.$
(ii) The proof is almost identical to that of (i), so we omit it here. ∎
Next, we shall prove Theorem 1.8, which is based on the following lemma.